# Doubly-charmed pentaquark states in a mass splitting model

### v1

@@ III.3 Spectra and decay widths of $ccnn\bar{s}$ states
$\frac{5}{2}^{-}$ $\left(\begin{array}[]{c}136.0\end{array}\right)$ $\left(\begin{array}[]{c}136.0\end{array}\right)$ [$\left[\begin{array}[]{c}\{1.00\}\end{array}\right]$ $\left(\begin{array}[]{c}4609.2\end{array}\right)$ $\left(\begin{array}[]{c}4658.7\end{array}\right)$ $\left(\begin{array}[]{c}4586.5\end{array}\right)$ $\left(\begin{array}[]{c}4850.2\end{array}\right)$
$\frac{3}{2}^{-}$ $\left(\begin{array}[]{cccc}-33.3&35.8&167.3&59.9\\ 35.8&80.8&-42.8&-5.2\\ 167.3&-42.8&73.2&-16.0\\ 59.9&-5.2&-16.0&36.9\end{array}\right)$ $\left(\begin{array}[]{c}200.2\\ 93.1\\ 46.9\\ -182.6\end{array}\right)$ [$\left[\begin{array}[]{cccc}\{-0.58,0.11,-0.79,-0.14\}\\ \{-0.25,-0.94,0.09,-0.20\}\\ \{0.07,-0.24,-0.25,0.93\}\\ \{0.77,-0.20,-0.55,-0.25\}\end{array}\right]$ $\left(\begin{array}[]{c}4673.4\\ 4566.3\\ 4520.1\\ 4290.6\end{array}\right)$ $\left(\begin{array}[]{c}4722.9\\ 4615.8\\ 4569.6\\ 4340.1\end{array}\right)$ $\left(\begin{array}[]{c}4650.8\\ 4543.6\\ 4497.4\\ 4268.0\end{array}\right)$ $\left(\begin{array}[]{c}4914.4\\ 4807.3\\ 4761.1\\ 4531.6\end{array}\right)$
$\frac{1}{2}^{-}$ $\left(\begin{array}[]{cccc}36.3&45.3&-105.8&-37.9\\ 45.3&-20.8&181.6&-85.6\\ -105.8&181.6&100.0&-16.0\\ -37.9&-85.6&-16.0&111.7\end{array}\right)$ $\left(\begin{array}[]{c}265.6\\ 155.6\\ 5.9\\ -199.9\end{array}\right)$ [$\left[\begin{array}[]{cccc}\{-0.18,0.55,0.75,-0.34\}\\ \{-0.58,-0.18,0.32,0.73\}\\ \{-0.70,-0.38,-0.16,-0.58\}\\ \{0.37,-0.73,0.56,-0.13\}\end{array}\right]$ $\left(\begin{array}[]{c}4738.8\\ 4628.8\\ 4479.1\\ 4273.3\end{array}\right)$ $\left(\begin{array}[]{c}4788.3\\ 4678.3\\ 4528.6\\ 4322.8\end{array}\right)$ $\left(\begin{array}[]{c}4716.1\\ 4606.1\\ 4456.4\\ 4250.7\end{array}\right)$ $\left(\begin{array}[]{c}4979.8\\ 4869.8\\ 4720.1\\ 4514.3\end{array}\right)$
@@
$\frac{5}{2}^{-}$ $\left(\begin{array}[]{c}126.4\end{array}\right)$ $\left(\begin{array}[]{c}126.4\end{array}\right)$ [$\left[\begin{array}[]{c}\{1.00\}\end{array}\right]$ $\left(\begin{array}[]{c}4599.6\end{array}\right)$ $\left(\begin{array}[]{c}4635.7\end{array}\right)$ $\left(\begin{array}[]{c}4576.9\end{array}\right)$ $\left(\begin{array}[]{c}4840.6\end{array}\right)$
$\frac{3}{2}^{-}$ $\left(\begin{array}[]{ccc}-162.9&149.4&167.3\\ 149.4&15.2&-42.8\\ 167.3&-42.8&-120.8\end{array}\right)$ $\left(\begin{array}[]{c}113.2\\ -17.7\\ -364.0\end{array}\right)$ [$\left[\begin{array}[]{ccc}\{0.58,0.77,0.28\}\\ \{-0.34,0.54,-0.77\}\\ \{0.74,-0.36,-0.57\}\end{array}\right]$ $\left(\begin{array}[]{c}4586.4\\ 4455.5\\ 4109.2\end{array}\right)$ $\left(\begin{array}[]{c}4622.5\\ 4491.6\\ 4145.3\end{array}\right)$ $\left(\begin{array}[]{c}4563.7\\ 4432.8\\ 4086.5\end{array}\right)$ $\left(\begin{array}[]{c}4827.4\\ 4696.5\\ 4350.2\end{array}\right)$
$\frac{1}{2}^{-}$ $\left(\begin{array}[]{cccc}-69.3&27.7&188.9&-105.8\\ 27.7&-134.4&0.0&65.6\\ 188.9&0.0&-158.4&181.6\\ -105.8&65.6&181.6&-174.4\end{array}\right)$ $\left(\begin{array}[]{c}91.1\\ -16.9\\ -142.3\\ -468.5\end{array}\right)$ [$\left[\begin{array}[]{cccc}\{-0.67,-0.15,-0.68,-0.24\}\\ \{-0.55,0.29,0.22,0.75\}\\ \{-0.14,-0.93,0.29,0.17\}\\ \{0.47,-0.16,-0.63,0.59\}\end{array}\right]$ $\left(\begin{array}[]{c}4564.3\\ 4456.3\\ 4330.9\\ 4004.7\end{array}\right)$ $\left(\begin{array}[]{c}4600.4\\ 4492.4\\ 4367.0\\ 4040.8\end{array}\right)$ $\left(\begin{array}[]{c}4541.6\\ 4433.6\\ 4308.3\\ 3982.1\end{array}\right)$ $\left(\begin{array}[]{c}4805.3\\ 4697.3\\ 4571.9\\ 4245.7\end{array}\right)$
@@
$\frac{5}{2}^{-}$ $\left(\begin{array}[]{c}102.7\end{array}\right)$ $\left(\begin{array}[]{c}102.7\end{array}\right)$ [$\left[\begin{array}[]{c}\{1.00\}\end{array}\right]$ $\left(\begin{array}[]{c}4666.5\end{array}\right)$ $\left(\begin{array}[]{c}4801.4\end{array}\right)$ $\left(\begin{array}[]{c}4643.8\end{array}\right)$ $\left(\begin{array}[]{c}4997.5\end{array}\right)$
$\frac{3}{2}^{-}$ $\left(\begin{array}[]{cccc}-66.0&36.1&167.3&59.0\\ 36.1&46.0&-40.4&-8.0\\ 167.3&-40.4&22.5&-17.2\\ 59.0&-8.0&-17.2&3.1\end{array}\right)$ $\left(\begin{array}[]{c}155.7\\ 57.1\\ 13.7\\ -220.9\end{array}\right)$ [$\left[\begin{array}[]{cccc}\{-0.61,0.10,-0.77,-0.15\}\\ \{-0.23,-0.96,0.09,-0.13\}\\ \{0.07,-0.17,-0.26,0.95\}\\ \{0.76,-0.20,-0.57,-0.25\}\end{array}\right]$ $\left(\begin{array}[]{c}4719.5\\ 4620.9\\ 4577.5\\ 4342.9\end{array}\right)$ $\left(\begin{array}[]{c}4854.5\\ 4755.9\\ 4712.4\\ 4477.9\end{array}\right)$ $\left(\begin{array}[]{c}4696.9\\ 4598.2\\ 4554.8\\ 4320.2\end{array}\right)$ $\left(\begin{array}[]{c}5050.5\\ 4951.9\\ 4908.5\\ 4673.9\end{array}\right)$
$\frac{1}{2}^{-}$ $\left(\begin{array}[]{cccc}0.8&45.6&-105.8&-37.3\\ 45.6&-55.2&184.0&-87.2\\ -105.8&184.0&48.9&-17.2\\ -37.3&-87.2&-17.2&77.9\end{array}\right)$ $\left(\begin{array}[]{c}225.4\\ 118.4\\ -29.2\\ -242.2\end{array}\right)$ [$\left[\begin{array}[]{cccc}\{0.16,-0.56,-0.72,0.37\}\\ \{-0.59,-0.14,0.35,0.71\}\\ \{-0.70,-0.39,-0.16,-0.58\}\\ \{0.37,-0.72,0.58,-0.12\}\end{array}\right]$ $\left(\begin{array}[]{c}4789.2\\ 4682.2\\ 4534.6\\ 4321.6\end{array}\right)$ $\left(\begin{array}[]{c}4924.2\\ 4817.1\\ 4669.5\\ 4456.6\end{array}\right)$ $\left(\begin{array}[]{c}4766.6\\ 4659.5\\ 4511.9\\ 4298.9\end{array}\right)$ $\left(\begin{array}[]{c}5120.2\\ 5013.2\\ 4865.6\\ 4652.6\end{array}\right)$
@@
$\frac{5}{2}^{-}$ $\left(\begin{array}[]{c}78.1\end{array}\right)$ $\left(\begin{array}[]{c}78.1\end{array}\right)$ [$\left[\begin{array}[]{c}\{1.00\}\end{array}\right]$ $\left(\begin{array}[]{c}4732.5\end{array}\right)$ $\left(\begin{array}[]{c}4879.6\end{array}\right)$ $\left(\begin{array}[]{c}4793.8\end{array}\right)$ $\left(\begin{array}[]{c}5153.5\end{array}\right)$
$\frac{3}{2}^{-}$ $\left(\begin{array}[]{cccc}-29.2&8.0&84.1&59.9\\ 8.0&33.7&-3.2&-7.6\\ 84.1&-3.2&22.4&-17.2\\ 59.9&-7.6&-17.2&15.5\end{array}\right)$ $\left(\begin{array}[]{c}92.6\\ 37.9\\ 30.1\\ -118.2\end{array}\right)$ [$\left[\begin{array}[]{cccc}\{0.64,0.01,0.68,0.35\}\\ \{0.03,-0.75,-0.32,0.58\}\\ \{0.11,0.66,-0.42,0.61\}\\ \{0.76,-0.07,-0.50,-0.41\}\end{array}\right]$ $\left(\begin{array}[]{c}4747.0\\ 4692.3\\ 4684.5\\ 4536.2\end{array}\right)$ $\left(\begin{array}[]{c}4894.1\\ 4839.4\\ 4831.5\\ 4683.2\end{array}\right)$ $\left(\begin{array}[]{c}4808.3\\ 4753.6\\ 4745.8\\ 4597.5\end{array}\right)$ $\left(\begin{array}[]{c}5168.0\\ 5113.3\\ 5105.5\\ 4957.2\end{array}\right)$
$\frac{1}{2}^{-}$ $\left(\begin{array}[]{cccc}0.8&10.2&-53.2&-37.9\\ 10.2&-30.7&109.6&-88.0\\ -53.2&109.6&49.2&-17.2\\ -37.9&-88.0&-17.2&53.1\end{array}\right)$ $\left(\begin{array}[]{c}167.8\\ 83.2\\ -44.5\\ -134.0\end{array}\right)$ [$\left[\begin{array}[]{cccc}\{0.05,-0.58,-0.63,0.52\}\\ \{-0.61,-0.04,0.50,0.61\}\\ \{0.77,0.18,0.32,0.52\}\\ \{0.17,-0.80,0.50,-0.29\}\end{array}\right]$ $\left(\begin{array}[]{c}4822.2\\ 4737.6\\ 4609.9\\ 4520.4\end{array}\right)$ $\left(\begin{array}[]{c}4969.2\\ 4884.7\\ 4756.9\\ 4667.4\end{array}\right)$ $\left(\begin{array}[]{c}4883.5\\ 4798.9\\ 4671.2\\ 4581.7\end{array}\right)$ $\left(\begin{array}[]{c}5243.2\\ 5158.6\\ 5030.9\\ 4941.4\end{array}\right)$
@@ III.4 Spectra and decay widths of $ccss\bar{n}$ and $ccss\bar{s}$ states
$\frac{5}{2}^{-}$ $\left(\begin{array}[]{cc}173.3&-32.5\\ -32.5&134.3\end{array}\right)$ $\left(\begin{array}[]{c}191.7\\ 115.8\end{array}\right)$ [$\left[\begin{array}[]{cc}\{-0.87,0.49\}\\ \{-0.49,-0.87\}\end{array}\right]$ $\left(\begin{array}[]{c}4664.9\\ 4589.0\end{array}\right)$ $\left(\begin{array}[]{c}4730.5\\ 4654.6\end{array}\right)$ $\left(\begin{array}[]{c}4642.2\\ 4566.4\end{array}\right)$ $\left(\begin{array}[]{c}4905.9\\ 4830.0\end{array}\right)$
$\frac{3}{2}^{-}$ $\left(\begin{array}[]{ccccccc}-209.7&46.7&191.0&-35.4&-59.0&217.3&0.0\\ 46.7&-71.7&-35.4&52.8&217.3&-23.6&59.0\\ 191.0&-35.4&41.3&-15.0&25.0&-64.0&-0.8\\ -35.4&52.8&-15.0&70.9&-64.0&10.0&-6.8\\ -59.0&217.3&25.0&-64.0&48.1&0.0&-16.6\\ 217.3&-23.6&-64.0&10.0&0.0&-70.7&-0.8\\ 0.0&59.0&-0.8&-6.8&-16.6&-0.8&12.7\end{array}\right)$ $\left(\begin{array}[]{c}223.2\\ 169.0\\ 83.4\\ 42.7\\ 19.7\\ -237.2\\ -479.9\end{array}\right)$ [$\left[\begin{array}[]{ccccccc}\{0.16,-0.57,0.11,0.09,-0.77,0.14,-0.10\}\\ \{0.52,0.05,0.71,-0.36,0.17,0.26,0.01\}\\ \{-0.26,-0.31,-0.08,-0.83,0.03,-0.33,-0.18\}\\ \{-0.21,-0.04,0.58,0.34,0.00,-0.69,-0.15\}\\ \{-0.05,0.04,0.07,-0.14,-0.21,-0.20,0.94\}\\ \{0.25,0.70,-0.11,-0.20,-0.53,-0.26,-0.21\}\\ \{-0.73,0.28,0.35,-0.10,-0.23,0.46,-0.04\}\end{array}\right]$ $\left(\begin{array}[]{c}4696.4\\ 4642.2\\ 4556.6\\ 4515.9\\ 4492.9\\ 4236.0\\ 3993.3\end{array}\right)$ $\left(\begin{array}[]{c}4762.0\\ 4707.8\\ 4622.2\\ 4581.5\\ 4558.5\\ 4301.6\\ 4058.9\end{array}\right)$ $\left(\begin{array}[]{c}4673.8\\ 4619.5\\ 4534.0\\ 4493.3\\ 4470.2\\ 4213.3\\ 3970.6\end{array}\right)$ $\left(\begin{array}[]{c}4937.4\\ 4883.2\\ 4797.6\\ 4756.9\\ 4733.9\\ 4477.0\\ 4234.3\end{array}\right)$
$\frac{1}{2}^{-}$ $\left(\begin{array}[]{cccccccc}-62.9&1.7&28.8&241.5&-44.8&37.3&-137.5&0.0\\ 1.7&18.7&1.5&-44.8&66.8&-137.5&14.9&-37.3\\ 28.8&1.5&-84.0&0.0&0.0&0.0&64.7&12.9\\ 241.5&-44.8&0.0&-188.5&32.5&-54.2&227.6&-0.8\\ -44.8&66.8&0.0&32.5&-52.7&227.6&-21.7&-86.0\\ 37.3&-137.5&0.0&-54.2&227.6&74.5&0.0&-16.6\\ -137.5&14.9&64.7&227.6&-21.7&0.0&-123.5&-0.8\\ 0.0&-37.3&12.9&-0.8&-86.0&-16.6&-0.8&109.9\end{array}\right)$ $\left(\begin{array}[]{c}285.9\\ 164.9\\ 132.8\\ 46.0\\ -0.7\\ -95.1\\ -267.9\\ -574.5\end{array}\right)$ [$\left[\begin{array}[]{cccccccc}\{0.01,0.20,0.03,0.07,-0.54,-0.75,0.08,0.29\}\\ \{0.38,-0.55,0.09,0.32,-0.18,0.25,0.08,0.59\}\\ \{-0.51,-0.16,-0.13,-0.60,-0.09,0.11,-0.29,0.48\}\\ \{0.56,0.01,-0.26,-0.16,-0.13,-0.05,-0.75,-0.14\}\\ \{-0.21,-0.69,-0.11,-0.05,-0.39,-0.12,0.07,-0.54\}\\ \{0.08,-0.03,0.94,-0.25,-0.12,0.02,-0.17,-0.11\}\\ \{0.10,0.37,-0.10,-0.15,-0.67,0.55,0.24,-0.09\}\\ \{0.48,-0.13,-0.09,-0.65,0.20,-0.18,0.50,0.02\}\end{array}\right]$ $\left(\begin{array}[]{c}4759.1\\ 4638.1\\ 4606.0\\ 4519.2\\ 4472.5\\ 4378.1\\ 4205.3\\ 3898.7\end{array}\right)$ $\left(\begin{array}[]{c}4824.7\\ 4703.7\\ 4671.6\\ 4584.8\\ 4538.1\\ 4443.7\\ 4270.9\\ 3964.3\end{array}\right)$ $\left(\begin{array}[]{c}4736.5\\ 4615.5\\ 4583.3\\ 4496.6\\ 4449.9\\ 4355.5\\ 4182.6\\ 3876.0\end{array}\right)$ $\left(\begin{array}[]{c}5000.1\\ 4879.1\\ 4847.0\\ 4760.2\\ 4713.5\\ 4619.1\\ 4446.3\\ 4139.7\end{array}\right)$

### v2

@@ III.3 Spectra and decay widths of $ccnn\bar{s}$ states
$\frac{5}{2}^{-}$ $\left(\begin{array}[]{c}136.0\end{array}\right)$ $\left(\begin{array}[]{c}136.0\end{array}\right)$ $\left[\begin{array}[]{c}\{1.00\}\end{array}\right]$ $\left(\begin{array}[]{c}4609.2\end{array}\right)$ $\left(\begin{array}[]{c}4658.7\end{array}\right)$ $\left(\begin{array}[]{c}4586.5\end{array}\right)$ $\left(\begin{array}[]{c}4850.2\end{array}\right)$
$\frac{3}{2}^{-}$ $\left(\begin{array}[]{cccc}-33.3&35.8&167.3&59.9\\ 35.8&80.8&-42.8&-5.2\\ 167.3&-42.8&73.2&-16.0\\ 59.9&-5.2&-16.0&36.9\end{array}\right)$ $\left(\begin{array}[]{c}200.2\\ 93.1\\ 46.9\\ -182.6\end{array}\right)$ $\left[\begin{array}[]{cccc}\{-0.58,0.11,-0.79,-0.14\}\\ \{-0.25,-0.94,0.09,-0.20\}\\ \{0.07,-0.24,-0.25,0.93\}\\ \{0.77,-0.20,-0.55,-0.25\}\end{array}\right]$ $\left(\begin{array}[]{c}4673.4\\ 4566.3\\ 4520.1\\ 4290.6\end{array}\right)$ $\left(\begin{array}[]{c}4722.9\\ 4615.8\\ 4569.6\\ 4340.1\end{array}\right)$ $\left(\begin{array}[]{c}4650.8\\ 4543.6\\ 4497.4\\ 4268.0\end{array}\right)$ $\left(\begin{array}[]{c}4914.4\\ 4807.3\\ 4761.1\\ 4531.6\end{array}\right)$
$\frac{1}{2}^{-}$ $\left(\begin{array}[]{cccc}36.3&45.3&-105.8&-37.9\\ 45.3&-20.8&181.6&-85.6\\ -105.8&181.6&100.0&-16.0\\ -37.9&-85.6&-16.0&111.7\end{array}\right)$ $\left(\begin{array}[]{c}265.6\\ 155.6\\ 5.9\\ -199.9\end{array}\right)$ $\left[\begin{array}[]{cccc}\{-0.18,0.55,0.75,-0.34\}\\ \{-0.58,-0.18,0.32,0.73\}\\ \{-0.70,-0.38,-0.16,-0.58\}\\ \{0.37,-0.73,0.56,-0.13\}\end{array}\right]$ $\left(\begin{array}[]{c}4738.8\\ 4628.8\\ 4479.1\\ 4273.3\end{array}\right)$ $\left(\begin{array}[]{c}4788.3\\ 4678.3\\ 4528.6\\ 4322.8\end{array}\right)$ $\left(\begin{array}[]{c}4716.1\\ 4606.1\\ 4456.4\\ 4250.7\end{array}\right)$ $\left(\begin{array}[]{c}4979.8\\ 4869.8\\ 4720.1\\ 4514.3\end{array}\right)$
@@
$\frac{5}{2}^{-}$ $\left(\begin{array}[]{c}126.4\end{array}\right)$ $\left(\begin{array}[]{c}126.4\end{array}\right)$ $\left[\begin{array}[]{c}\{1.00\}\end{array}\right]$ $\left(\begin{array}[]{c}4599.6\end{array}\right)$ $\left(\begin{array}[]{c}4635.7\end{array}\right)$ $\left(\begin{array}[]{c}4576.9\end{array}\right)$ $\left(\begin{array}[]{c}4840.6\end{array}\right)$
$\frac{3}{2}^{-}$ $\left(\begin{array}[]{ccc}-162.9&149.4&167.3\\ 149.4&15.2&-42.8\\ 167.3&-42.8&-120.8\end{array}\right)$ $\left(\begin{array}[]{c}113.2\\ -17.7\\ -364.0\end{array}\right)$ $\left[\begin{array}[]{ccc}\{0.58,0.77,0.28\}\\ \{-0.34,0.54,-0.77\}\\ \{0.74,-0.36,-0.57\}\end{array}\right]$ $\left(\begin{array}[]{c}4586.4\\ 4455.5\\ 4109.2\end{array}\right)$ $\left(\begin{array}[]{c}4622.5\\ 4491.6\\ 4145.3\end{array}\right)$ $\left(\begin{array}[]{c}4563.7\\ 4432.8\\ 4086.5\end{array}\right)$ $\left(\begin{array}[]{c}4827.4\\ 4696.5\\ 4350.2\end{array}\right)$
$\frac{1}{2}^{-}$ $\left(\begin{array}[]{cccc}-69.3&27.7&188.9&-105.8\\ 27.7&-134.4&0.0&65.6\\ 188.9&0.0&-158.4&181.6\\ -105.8&65.6&181.6&-174.4\end{array}\right)$ $\left(\begin{array}[]{c}91.1\\ -16.9\\ -142.3\\ -468.5\end{array}\right)$ $\left[\begin{array}[]{cccc}\{-0.67,-0.15,-0.68,-0.24\}\\ \{-0.55,0.29,0.22,0.75\}\\ \{-0.14,-0.93,0.29,0.17\}\\ \{0.47,-0.16,-0.63,0.59\}\end{array}\right]$ $\left(\begin{array}[]{c}4564.3\\ 4456.3\\ 4330.9\\ 4004.7\end{array}\right)$ $\left(\begin{array}[]{c}4600.4\\ 4492.4\\ 4367.0\\ 4040.8\end{array}\right)$ $\left(\begin{array}[]{c}4541.6\\ 4433.6\\ 4308.3\\ 3982.1\end{array}\right)$ $\left(\begin{array}[]{c}4805.3\\ 4697.3\\ 4571.9\\ 4245.7\end{array}\right)$
@@
$\frac{5}{2}^{-}$ $\left(\begin{array}[]{c}102.7\end{array}\right)$ $\left(\begin{array}[]{c}102.7\end{array}\right)$ $\left[\begin{array}[]{c}\{1.00\}\end{array}\right]$ $\left(\begin{array}[]{c}4666.5\end{array}\right)$ $\left(\begin{array}[]{c}4801.4\end{array}\right)$ $\left(\begin{array}[]{c}4643.8\end{array}\right)$ $\left(\begin{array}[]{c}4997.5\end{array}\right)$
$\frac{3}{2}^{-}$ $\left(\begin{array}[]{cccc}-66.0&36.1&167.3&59.0\\ 36.1&46.0&-40.4&-8.0\\ 167.3&-40.4&22.5&-17.2\\ 59.0&-8.0&-17.2&3.1\end{array}\right)$ $\left(\begin{array}[]{c}155.7\\ 57.1\\ 13.7\\ -220.9\end{array}\right)$ $\left[\begin{array}[]{cccc}\{-0.61,0.10,-0.77,-0.15\}\\ \{-0.23,-0.96,0.09,-0.13\}\\ \{0.07,-0.17,-0.26,0.95\}\\ \{0.76,-0.20,-0.57,-0.25\}\end{array}\right]$ $\left(\begin{array}[]{c}4719.5\\ 4620.9\\ 4577.5\\ 4342.9\end{array}\right)$ $\left(\begin{array}[]{c}4854.5\\ 4755.9\\ 4712.4\\ 4477.9\end{array}\right)$ $\left(\begin{array}[]{c}4696.9\\ 4598.2\\ 4554.8\\ 4320.2\end{array}\right)$ $\left(\begin{array}[]{c}5050.5\\ 4951.9\\ 4908.5\\ 4673.9\end{array}\right)$
$\frac{1}{2}^{-}$ $\left(\begin{array}[]{cccc}0.8&45.6&-105.8&-37.3\\ 45.6&-55.2&184.0&-87.2\\ -105.8&184.0&48.9&-17.2\\ -37.3&-87.2&-17.2&77.9\end{array}\right)$ $\left(\begin{array}[]{c}225.4\\ 118.4\\ -29.2\\ -242.2\end{array}\right)$ $\left[\begin{array}[]{cccc}\{0.16,-0.56,-0.72,0.37\}\\ \{-0.59,-0.14,0.35,0.71\}\\ \{-0.70,-0.39,-0.16,-0.58\}\\ \{0.37,-0.72,0.58,-0.12\}\end{array}\right]$ $\left(\begin{array}[]{c}4789.2\\ 4682.2\\ 4534.6\\ 4321.6\end{array}\right)$ $\left(\begin{array}[]{c}4924.2\\ 4817.1\\ 4669.5\\ 4456.6\end{array}\right)$ $\left(\begin{array}[]{c}4766.6\\ 4659.5\\ 4511.9\\ 4298.9\end{array}\right)$ $\left(\begin{array}[]{c}5120.2\\ 5013.2\\ 4865.6\\ 4652.6\end{array}\right)$
@@
$\frac{5}{2}^{-}$ $\left(\begin{array}[]{c}78.1\end{array}\right)$ $\left(\begin{array}[]{c}78.1\end{array}\right)$ $\left[\begin{array}[]{c}\{1.00\}\end{array}\right]$ $\left(\begin{array}[]{c}4732.5\end{array}\right)$ $\left(\begin{array}[]{c}4879.6\end{array}\right)$ $\left(\begin{array}[]{c}4793.8\end{array}\right)$ $\left(\begin{array}[]{c}5153.5\end{array}\right)$
$\frac{3}{2}^{-}$ $\left(\begin{array}[]{cccc}-29.2&8.0&84.1&59.9\\ 8.0&33.7&-3.2&-7.6\\ 84.1&-3.2&22.4&-17.2\\ 59.9&-7.6&-17.2&15.5\end{array}\right)$ $\left(\begin{array}[]{c}92.6\\ 37.9\\ 30.1\\ -118.2\end{array}\right)$ $\left[\begin{array}[]{cccc}\{0.64,0.01,0.68,0.35\}\\ \{0.03,-0.75,-0.32,0.58\}\\ \{0.11,0.66,-0.42,0.61\}\\ \{0.76,-0.07,-0.50,-0.41\}\end{array}\right]$ $\left(\begin{array}[]{c}4747.0\\ 4692.3\\ 4684.5\\ 4536.2\end{array}\right)$ $\left(\begin{array}[]{c}4894.1\\ 4839.4\\ 4831.5\\ 4683.2\end{array}\right)$ $\left(\begin{array}[]{c}4808.3\\ 4753.6\\ 4745.8\\ 4597.5\end{array}\right)$ $\left(\begin{array}[]{c}5168.0\\ 5113.3\\ 5105.5\\ 4957.2\end{array}\right)$
$\frac{1}{2}^{-}$ $\left(\begin{array}[]{cccc}0.8&10.2&-53.2&-37.9\\ 10.2&-30.7&109.6&-88.0\\ -53.2&109.6&49.2&-17.2\\ -37.9&-88.0&-17.2&53.1\end{array}\right)$ $\left(\begin{array}[]{c}167.8\\ 83.2\\ -44.5\\ -134.0\end{array}\right)$ $\left[\begin{array}[]{cccc}\{0.05,-0.58,-0.63,0.52\}\\ \{-0.61,-0.04,0.50,0.61\}\\ \{0.77,0.18,0.32,0.52\}\\ \{0.17,-0.80,0.50,-0.29\}\end{array}\right]$ $\left(\begin{array}[]{c}4822.2\\ 4737.6\\ 4609.9\\ 4520.4\end{array}\right)$ $\left(\begin{array}[]{c}4969.2\\ 4884.7\\ 4756.9\\ 4667.4\end{array}\right)$ $\left(\begin{array}[]{c}4883.5\\ 4798.9\\ 4671.2\\ 4581.7\end{array}\right)$ $\left(\begin{array}[]{c}5243.2\\ 5158.6\\ 5030.9\\ 4941.4\end{array}\right)$
@@ III.4 Spectra and decay widths of $ccss\bar{n}$ and $ccss\bar{s}$ states
$\frac{5}{2}^{-}$ $\left(\begin{array}[]{cc}173.3&-32.5\\ -32.5&134.3\end{array}\right)$ $\left(\begin{array}[]{c}191.7\\ 115.8\end{array}\right)$ $\left[\begin{array}[]{cc}\{-0.87,0.49\}\\ \{-0.49,-0.87\}\end{array}\right]$ $\left(\begin{array}[]{c}4664.9\\ 4589.0\end{array}\right)$ $\left(\begin{array}[]{c}4730.5\\ 4654.6\end{array}\right)$ $\left(\begin{array}[]{c}4642.2\\ 4566.4\end{array}\right)$ $\left(\begin{array}[]{c}4905.9\\ 4830.0\end{array}\right)$
$\frac{3}{2}^{-}$ $\left(\begin{array}[]{ccccccc}-209.7&46.7&191.0&-35.4&-59.0&217.3&0.0\\ 46.7&-71.7&-35.4&52.8&217.3&-23.6&59.0\\ 191.0&-35.4&41.3&-15.0&25.0&-64.0&-0.8\\ -35.4&52.8&-15.0&70.9&-64.0&10.0&-6.8\\ -59.0&217.3&25.0&-64.0&48.1&0.0&-16.6\\ 217.3&-23.6&-64.0&10.0&0.0&-70.7&-0.8\\ 0.0&59.0&-0.8&-6.8&-16.6&-0.8&12.7\end{array}\right)$ $\left(\begin{array}[]{c}223.2\\ 169.0\\ 83.4\\ 42.7\\ 19.7\\ -237.2\\ -479.9\end{array}\right)$ $\left[\begin{array}[]{ccccccc}\{0.16,-0.57,0.11,0.09,-0.77,0.14,-0.10\}\\ \{0.52,0.05,0.71,-0.36,0.17,0.26,0.01\}\\ \{-0.26,-0.31,-0.08,-0.83,0.03,-0.33,-0.18\}\\ \{-0.21,-0.04,0.58,0.34,0.00,-0.69,-0.15\}\\ \{-0.05,0.04,0.07,-0.14,-0.21,-0.20,0.94\}\\ \{0.25,0.70,-0.11,-0.20,-0.53,-0.26,-0.21\}\\ \{-0.73,0.28,0.35,-0.10,-0.23,0.46,-0.04\}\end{array}\right]$ $\left(\begin{array}[]{c}4696.4\\ 4642.2\\ 4556.6\\ 4515.9\\ 4492.9\\ 4236.0\\ 3993.3\end{array}\right)$ $\left(\begin{array}[]{c}4762.0\\ 4707.8\\ 4622.2\\ 4581.5\\ 4558.5\\ 4301.6\\ 4058.9\end{array}\right)$ $\left(\begin{array}[]{c}4673.8\\ 4619.5\\ 4534.0\\ 4493.3\\ 4470.2\\ 4213.3\\ 3970.6\end{array}\right)$ $\left(\begin{array}[]{c}4937.4\\ 4883.2\\ 4797.6\\ 4756.9\\ 4733.9\\ 4477.0\\ 4234.3\end{array}\right)$
$\frac{1}{2}^{-}$ $\left(\begin{array}[]{cccccccc}-62.9&1.7&28.8&241.5&-44.8&37.3&-137.5&0.0\\ 1.7&18.7&1.5&-44.8&66.8&-137.5&14.9&-37.3\\ 28.8&1.5&-84.0&0.0&0.0&0.0&64.7&12.9\\ 241.5&-44.8&0.0&-188.5&32.5&-54.2&227.6&-0.8\\ -44.8&66.8&0.0&32.5&-52.7&227.6&-21.7&-86.0\\ 37.3&-137.5&0.0&-54.2&227.6&74.5&0.0&-16.6\\ -137.5&14.9&64.7&227.6&-21.7&0.0&-123.5&-0.8\\ 0.0&-37.3&12.9&-0.8&-86.0&-16.6&-0.8&109.9\end{array}\right)$ $\left(\begin{array}[]{c}285.9\\ 164.9\\ 132.8\\ 46.0\\ -0.7\\ -95.1\\ -267.9\\ -574.5\end{array}\right)$ $\left[\begin{array}[]{cccccccc}\{0.01,0.20,0.03,0.07,-0.54,-0.75,0.08,0.29\}\\ \{0.38,-0.55,0.09,0.32,-0.18,0.25,0.08,0.59\}\\ \{-0.51,-0.16,-0.13,-0.60,-0.09,0.11,-0.29,0.48\}\\ \{0.56,0.01,-0.26,-0.16,-0.13,-0.05,-0.75,-0.14\}\\ \{-0.21,-0.69,-0.11,-0.05,-0.39,-0.12,0.07,-0.54\}\\ \{0.08,-0.03,0.94,-0.25,-0.12,0.02,-0.17,-0.11\}\\ \{0.10,0.37,-0.10,-0.15,-0.67,0.55,0.24,-0.09\}\\ \{0.48,-0.13,-0.09,-0.65,0.20,-0.18,0.50,0.02\}\end{array}\right]$ $\left(\begin{array}[]{c}4759.1\\ 4638.1\\ 4606.0\\ 4519.2\\ 4472.5\\ 4378.1\\ 4205.3\\ 3898.7\end{array}\right)$ $\left(\begin{array}[]{c}4824.7\\ 4703.7\\ 4671.6\\ 4584.8\\ 4538.1\\ 4443.7\\ 4270.9\\ 3964.3\end{array}\right)$ $\left(\begin{array}[]{c}4736.5\\ 4615.5\\ 4583.3\\ 4496.6\\ 4449.9\\ 4355.5\\ 4182.6\\ 3876.0\end{array}\right)$ $\left(\begin{array}[]{c}5000.1\\ 4879.1\\ 4847.0\\ 4760.2\\ 4713.5\\ 4619.1\\ 4446.3\\ 4139.7\end{array}\right)$